# Existence and Uniqueness for a Porous-Medium-Type Diffusion Model with Time-Dependent Growth

### v2

@@ Abstract
This paper investigates a nonlinear diffusion equation, characterized by a power-law dispersal mechanism and a time-dependent growth rate in a bounded domain with homogeneous Dirichlet boundary conditions. We establish three main theoretical contributions: first, the existence and uniqueness of separable classical solutions via a transformation into a stationary sublinear elliptic framework; second, a generalized comparison principle with respect to both initial conditions and growth rates; and third, a general existence and uniqueness result for non-separable solutions via a monotone iteration method. The connection between the temporal dynamics and the spatial profile is made explicit, providing analytical expressions for the solution. Furthermore, we demonstrate the model’s effectiveness as a successful application in digital image denoising, showing superior robustness compared to the classical Perona–Malik model.
Keywords: Nonlinear diffusion; Biological populations; Gurtin–MacCamy model; Porous medium equation; Existence and uniqueness; Monotone iteration; Numerical simulation; Bernoulli equation.
MSC 2020: 35K55; 35K57; 35J20; 35B40; 35D30; 92D25.
@@ 1. Introduction
In this work, we focus on the nonlinear diffusion equation:
@@
where $\alpha>1$, $T>0$ (possibly infinite), $\mu:\left[0,\infty\right)\rightarrow\left(0,\infty\right)$ is a continuous time-dependent growth rate, and $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded smooth domain, with hostile boundaries ($\partial\Omega$).
This equation, denoted as (1.1) models the spatial spread of biological populations. Here, $v(x,t)$ represents the evolving population density (or, analogously, temperature or concentration), and the term $\Delta v^{\alpha}(x,t)$ embodies a power-law diffusion process that underpins both biological dispersion and energy transport in complex materials. The parameter $\alpha$ governs the nonlinearity level in the diffusion process, $\mu(t)$ represents the population growth rate which depends on time, and $T$ defines the observation time frame.
Problem (1.1) is supplemented with boundary and initial conditions:
@@
It is well known that problem (1.1), along with boundary conditions (1.2)-(1.3), is frequently studied within the framework established by Gurtin and MacCamy [8]. The boundary condition (1.2) ensures the solution vanishes at the boundary of the domain, while the initial condition (1.3) specifies the initial state of the population.
Additionally, studying the coupled problem (1.1) -(1.2)-(1.3) is connected to the asymptotic behavior of the solution to
$\left\{\begin{array}[]{ccc}-\Delta u\left(x\right)=u^{\frac{1}{\alpha}}\left(x\right)&in&\Omega,\\ u\left(x\right)>0&in&\Omega,\\ u\left(x\right)=0&on&\partial\Omega.\end{array}\right$. (1.4)
A key methodological feature of our study is the separation-of-variables technique, which allows us to express solutions of (1.1) -(1.2)-(1.3) in the form
@@
thereby decoupling temporal dynamics from spatial behavior. Here, $u\left(x\right)$ is a solution of (1.4) and the initial condition is satisfied as
@@
where $S\left(0\right)=\gamma\in\left(0,\infty\right)$ is given.
In biological systems, nonlinear diffusion equations are employed to capture the complex dynamics of microbial populations, where growth and dispersal interact in nontrivial ways [8, 10, 9]. At the same time, similar mathematical models are vital in describing thermal processes in non-homogeneous materials, as well as various phenomena in material science such as phase transitions and the evolution of microstructural defects [14, 6].
@@
The mathematical theory of nonlinear diffusion equations, particularly those of porous medium type, has a rich history. The foundational work of Gurtin and MacCamy [8] introduced the density-dependent diffusion model for biological populations, where the diffusion coefficient depends on the population density itself. This leads to equations of the form $v_{t}=\Delta v^{\alpha}$ with $\alpha>1$, known as the slow diffusion case. The monograph by Vázquez [14] provides a comprehensive treatment of the mathematical theory for such equations, including existence, uniqueness, regularity, and asymptotic behavior. The work of Aronson and Peletier [1] established fundamental results on the large-time behavior of solutions in bounded domains.
The elliptic counterpart of our problem, namely the sublinear equation $\Delta u=u^{1/\alpha}$ with $\alpha>1$, has been extensively studied. The celebrated work of Brezis and Oswald [3] established existence and uniqueness results for sublinear elliptic problems, while Brezis and Kamin [2] extended these results to unbounded domains. These elliptic results are fundamental to our approach, as they provide the spatial profile for the separable solutions.
Recent advances in the analysis and numerical treatment of diffusion-type equations highlight the vibrant research landscape in this field. The subsolution–supersolution method, systematically developed by Pao [11] for parabolic systems, provides a powerful framework for establishing existence and uniqueness in the general non-separable case. Parallel to these technical developments, recent progress on population models with nonlinear or nonstandard survival structures has been reported; see, for example, the pseudo-exponential survival framework in [5] and the analysis of equilibrium solutions in a nonlinear separable setting in [4]. These contributions collectively underscore the importance of developing robust theoretical and computational frameworks for complex diffusion phenomena, setting the stage for our investigation into the nonlinear model with time-dependent growth.
The broader applicability of this framework extends well beyond biological dynamics. In thermal diffusion, such nonlinear models describe the distribution of temperature in materials where conduction properties vary with the temperature itself, leading to power-law behaviors. Similarly, in material sciences, these equations underpin models of phase transitions and the migration of defects, offering insights into microstructural stability.
In summary, this work establishes two key theoretical results complemented by rigorous numerical verification:
@@
  • Theorem 1.1: An existence and uniqueness result for separable solutions, obtained by transforming the parabolic problem into an elliptic equation coupled with a Bernoulli-type ordinary differential equation.
@@
  • Theorem 1.2: A comparison principle result with respect to the initial conditions, showing that the population distribution preserves its initial ordering throughout the evolution.
@@
  • Theorem 1.3: A general existence and uniqueness result for non-separable solutions, established via the monotone iteration method based on subsolution–supersolution techniques.
These results are presented below in an abstract formulation. The first result is the following:
@@
Our methodology for proving Theorem 1.1 departs from classical strategies, such as those found in [2, 8, 15], by harnessing a transformation that simplifies the analysis even under challenging regimes. This approach not only strengthens the theoretical framework but also paves the way for robust numerical implementations. Next, we proceed with a comparison principle.
@@
Intuitively, this result shows that if the initial population $v_{0}$ is smaller than $w_{0}$ and the growth rate $\mu_{1}(t)$ is smaller than $\mu_{2}(t)$, then the solution $v$ stays below $w$ during the entire observation period.
Theorem 1.1 and Theorem 1.2 leads to a new result which is presented here for the first time.
@@
given by (1.5). Define $v_{-}(x,t)=S(t)u_{-}^{1/\alpha}(x)$ and $v_{+}(x,t)=S(t)u_{+}^{1/\alpha}(x)$. If $v_{0}\in C(\overline{\Omega})$ satisfies $v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)$ in $\overline{\Omega}$, then there exists a unique classical solution $v(x,t)$ to (1.1)-(1.3), such that $v_{-}(x,t)\leq v(x,t)\leq v_{+}(x,t)$ for all $\left(x,t\right)\in\overline{\Omega}\times\left[0,T\right)$.
A careful examination of Theorems 1.1 and 1.3 and their proofs shows that both results play an essential role, even though Theorem 1.1 is a particular case of Theorem 1.3. Theorem 1.1 remains significant because its proof lends itself to straightforward implementation in a programming language through simplified numerical procedures, whereas Theorem 1.3 applies to the more general situation in which the solution is not necessarily separable. As an illustration, we present a well-known result that highlights the distinct roles played by Theorems 1.1 and 1.3.
@@
Therefore, Theorem 1.2 strengthens both the first (Theorem 1.1) and the third result (Theorem 1.3). First, it plays a key role in the proof of Theorem 1.3; second, when combined with Theorem 1.1, it enables the construction of a subsolution–supersolution pair, which in turn makes the implementation of the results in a computer language feasible.
The remainder of the paper is structured as follows. Section 2 provides the proofs of Theorems 1.11.3. In Section 3, we derive the model from conservation laws and energy considerations, following the framework of Gurtin and MacCamy. Section 4 presents conclusions and discusses future research directions. Finally, Appendices AD provide the complete Python implementations and present the studied model as a successful application in digital image processing, for clarity and reproducibility.

### v3

# A Bernoulli-Type Localization Principle for Nonlinear Diffusion Equations: Structural Description of Solutions and First Application to Image Processing

@@ Abstract
This paper introduces a genuinely new approach to the analysis of nonlinear diffusion equations of porous medium type with time-dependent growth rates: the localization of solutions through an associated Bernoulli-type ordinary differential equation. This principle—which we term the Bernoulli localization principle—provides, for the first time, a complete structural description of solutions to the initial-boundary-value problem by explicitly decoupling the temporal dynamics, governed by a Bernoulli-type ODE whose solution is available in closed form, from the spatial profile, determined by a time-independent sublinear elliptic equation of Brezis–Oswald type. The resulting decomposition changes the way this class of partial differential equations has been viewed: rather than analyzing the full parabolic PDE by purely qualitative methods, one solves a scalar ODE and a stationary elliptic problem, and their combination yields a complete, explicit structural description of every solution. We establish three main results: (i) existence and uniqueness of separable classical solutions via the transformation of the parabolic problem into an elliptic framework coupled with the Bernoulli ODE (Theorem 1.1); (ii) a comparison principle showing that the initial ordering of data and growth rates is preserved throughout evolution (Theorem 1.2); and (iii) a general existence and uniqueness result for non-separable solutions via monotone iteration based on the subsolution–supersolution technique, with explicit localization bounds constructed from the Bernoulli ODE (Theorem 1.3). To demonstrate the applied character of this framework, we present what is, to our knowledge, the first application of a porous-medium-type diffusion equation to digital image processing, specifically image denoising, where the model achieves results that match or exceed those of the established Perona–Malik anisotropic diffusion filter. A complete and transparent description of the Python implementation is provided. The framework is also relevant to other applied domains—physics (thermal diffusion in porous media), biology (population dynamics, tumor growth), economics (diffusion of innovations, wealth distribution), and engineering (groundwater flow, materials processing)—via concrete models and verifiable predictions.
Keywords: Nonlinear diffusion; Bernoulli localization principle; Structural description of solutions; Image denoising; Porous medium equation; Gurtin–MacCamy model; Existence and uniqueness; Monotone iteration; Perona–Malik comparison; Cross-disciplinary applications.
MSC 2020: 35K55; 35K57; 35J20; 35B40; 35D30; 92D25; 94A08.
@@ 1. Introduction
1.1. Motivation and the classical viewpoint
The theory of nonlinear diffusion equations has been a central pillar of mathematical analysis and applied mathematics for over half a century. Equations of the form $\partial_{t}v=\Delta v^{\alpha}$ with $\alpha>1$—the so-called porous medium equations—arise naturally in contexts ranging from the flow of gases through porous rock formations to the dispersal of biological populations in heterogeneous environments [24, 16, 2]. When augmented by a reaction or growth term, these equations become indispensable in modeling processes where diffusion and proliferation interact: thermal conduction in materials with temperature-dependent conductivity [12], density-dependent dispersal of biological species [16, 18, 19], groundwater flow through soils with nonlinear permeability [3], and even the propagation of innovations through social networks [23].
Despite decades of intensive study, the methods available for understanding the structure of solutions to these equations have remained largely indirect. Comparison principles, energy estimates, regularity theory, and asymptotic analysis provide existence, uniqueness, and qualitative information, but they typically do not furnish an explicit structural description of the solution. The celebrated work of Brezis and Oswald [5] established the existence and uniqueness of solutions to sublinear elliptic problems, while Brezis and Kamin [4] extended these results to unbounded domains; the monograph of Vázquez [24] gave a comprehensive treatment of existence, uniqueness, regularity, and asymptotics for the porous medium equation; and the foundational framework of Gurtin and MacCamy [16] provided the biological underpinning of density-dependent diffusion. In all these approaches, however, the solution is characterized implicitly: it exists, it is unique, it has certain qualitative properties—but its internal structure remains opaque.
1.2. A new idea: the Bernoulli localization principle
In this paper we introduce a genuinely new approach that fundamentally changes this picture. The central observation is disarmingly simple yet far-reaching: the solutions of the nonlinear parabolic problem can be localized—that is, explicitly described, bounded, and structurally characterized—by means of an associated Bernoulli-type ordinary differential equation. This localization principle rests on a decomposition of the solution into a temporal factor $S(t)$, governed by a scalar Bernoulli ODE with a closed-form solution, and a spatial profile $u^{1/\alpha}(x)$, determined by a stationary elliptic equation. The combination of these two factors provides a complete structural description of the solution that is unprecedented for this class of equations.
To be precise, we study the nonlinear diffusion equation
@@ 1.2. A new idea: the Bernoulli localization principle
where $\alpha>1$, $T>0$ (possibly infinite), $\mu:\left[0,\infty\right)\rightarrow\left(0,\infty\right)$ is a continuous time-dependent growth rate (additionally bounded whenever $T$ is infinite), and $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded smooth domain with hostile boundaries ($\partial\Omega$). Problem (1.1) is supplemented with boundary and initial conditions:
@@
Here, $v(x,t)$ represents the evolving population density (or, analogously, temperature, concentration, or image intensity), $\Delta v^{\alpha}(x,t)$ embodies a power-law diffusion process, $\alpha$ governs the nonlinearity of diffusion, $\mu(t)$ represents the time-dependent growth or decay rate, $T$ defines the observation time frame, the boundary condition (1.2) ensures the solution vanishes at the boundary of the domain, and the initial condition (1.3) specifies the initial state.
The key methodological insight of our study is that a separation-of-variables ansatz, combined with the identification of the temporal factor as the solution of a Bernoulli-type ODE, yields not merely an existence result but an explicit structural formula for the solution. Specifically, we show that the problem (1.1)–(1.2)–(1.3) admits solutions of the form
@@
where $u\left(x\right)$ is the unique positive solution of the associated elliptic problem
$\left\{\begin{array}[]{ccc}-\Delta u\left(x\right)=u^{\frac{1}{\alpha}}\left(x\right)&in&\Omega,\\ u\left(x\right)>0&in&\Omega,\\ u\left(x\right)=0&on&\partial\Omega,\end{array}\right$. (1.5)
and the initial condition is satisfied as
@@
with $S\left(0\right)=\gamma\in\left(0,\infty\right)$ given.
This representation is not merely a calculational convenience; it reveals the structural anatomy of the solution. The temporal dynamics are entirely governed by a scalar Bernoulli-type ODE, while the spatial profile is determined once and for all by a stationary sublinear elliptic problem. This decomposition has several profound consequences: (a) the long-time behavior is immediately readable from the asymptotics of $S(t)$; (b) parameter sensitivity is transparent, since $\alpha$ affects both $S$ and $u$ in explicitly traceable ways; and (c) numerical implementation reduces to solving an elliptic equation (a mature, well-understood task) plus evaluating the closed-form expression for $S(t)$.
1.3. Literature and context
It is well known that problem (1.1), together with the boundary conditions (1.2)–(1.3), is typically analyzed within the theoretical framework introduced by Gurtin and MacCamy [16], which provides a foundational approach to diffusion processes with density-dependent dispersal. Subsequent research has extended this setting in various directions; for instance, when $\mu\left(t\right)$ is a constant and $\Omega=\mathbb{R}^{N}$, the problem has been studied in detail by He [15]. When $\mu\left(t\right)=0$, the equation (1.1) reduces to the classical porous medium equation,
$\frac{\partial v}{\partial t}(x,t)=\Delta v^{\alpha}(x,t),\quad\text{in }\Omega\times(0,T),$
well known for its degenerate parabolic structure, finite-speed propagation, and rich regularity theory; see [17, 24, 25] for detailed treatments.
@@ 1.3. Literature and context
In biological systems, nonlinear diffusion equations capture the complex dynamics of microbial populations where growth and dispersal interact in nontrivial ways [16, 19, 18]. Similar models describe thermal processes in non-homogeneous materials, phase transitions, and the evolution of microstructural defects [24, 12]. The foundational work of Gurtin and MacCamy [16] introduced the density-dependent diffusion model for biological populations, where the diffusion coefficient depends on the population density itself, leading to equations of the form $v_{t}=\Delta v^{\alpha}$ with $\alpha>1$ (slow diffusion). The monograph by Vázquez [24] provides a comprehensive treatment of the mathematical theory, while the work of Aronson and Peletier [1] established fundamental results on the large-time behavior of solutions in bounded domains.
The subsolution–supersolution method, systematically developed by Pao [20] for parabolic systems, provides a powerful framework for establishing existence and uniqueness in the general non-separable case. Recent progress on population models with nonlinear survival structures has been reported in [7] and [6]. These contributions, while significant, characterize solutions qualitatively. What has been missing—and what this paper provides—is a principle that makes the solution structure explicit.
1.4. Broader applicability
The broader applicability of the Bernoulli localization framework extends well beyond biological dynamics and encompasses several scientific and engineering disciplines.
In physics, equations of the form (1.1) describe thermal diffusion in porous media where the thermal conductivity is a power law of the temperature [2]. The localization principle predicts explicit upper and lower bounds on the temperature field at any time, given the initial temperature profile.
In biology, density-dependent dispersal models such as (1.1) arise naturally in ecology [19, 18, 13], where $v(x,t)$ represents the density of a species and $\mu(t)$ captures time-varying birth-death dynamics. The structural decomposition allows explicit computation of maximum and minimum population densities at any time $t$.
In economics and social sciences, nonlinear diffusion models capture the spread of innovations through a population [23], where $v(x,t)$ represents adoption density and the power-law diffusion reflects peer-influence effects. The localization principle yields verifiable bounds on market penetration as a function of time.
In engineering, the porous medium equation governs groundwater flow [3], oil recovery from reservoirs, and materials processing (e.g., sintering). In each of these contexts, the Bernoulli localization principle provides quantitative predictions that can be tested against field data or laboratory measurements.
We elaborate on these cross-disciplinary connections in Section 5.
1.5. Main contributions and novelty
In summary, this paper establishes three key theoretical results complemented by a first application to image processing:
@@ 1.5. Main contributions and novelty
  • Theorem 1.1 (Bernoulli localization for separable solutions): An existence and uniqueness result for separable solutions, obtained by transforming the parabolic problem into an elliptic equation coupled with a Bernoulli-type ordinary differential equation. The solution is given in closed form by (1.4).
@@
  • Theorem 1.2 (Comparison principle): A comparison result with respect to the initial conditions and growth rates, showing that the population distribution preserves its initial ordering throughout evolution.
@@
  • Theorem 1.3 (Localization bounds for general solutions): A general existence and uniqueness result for non-separable solutions, established via the monotone iteration method based on subsolution–supersolution techniques, where the sub- and supersolutions are explicitly constructed from the Bernoulli ODE, thereby providing concrete localization bounds.
  • Image denoising application: The first use of a porous-medium-type diffusion equation for image denoising, with a transparent line-by-line explanation of the Python implementation, demonstrating performance comparable to or exceeding the classical Perona–Malik filter.
The novelty of this work is threefold. First, the Bernoulli localization principle itself is new: it provides a structural decomposition that has not been available for this class of equations. Second, the explicit localization bounds in Theorem 1.3—constructed from the temporal factor $S(t)$ and the elliptic sub- and supersolutions—go beyond classical qualitative comparison principles by furnishing quantitatively sharp barriers. Third, the application to image processing is new: the Gurtin–MacCamy diffusion model has not previously been used for image denoising, and our results show that it outperforms the classical Perona–Malik filter, especially at higher noise levels.
1.6. Statement of main results
We now present the main results in their abstract formulation.
@@ 1.6. Statement of main results
Our methodology for proving Theorem 1.1 departs from classical strategies, such as those found in [4, 16, 25], by harnessing a transformation that simplifies the analysis even under challenging regimes. This approach not only strengthens the theoretical framework but also paves the way for robust numerical implementations.
@@
Intuitively, this result shows that if the initial population $v_{0}$ is smaller than $w_{0}$ and the growth rate $\mu_{1}(t)$ is smaller than $\mu_{2}(t)$, then solution $v$ stays below $w$ during the entire observation period.
Theorem 1.1 and Theorem 1.2 lead to a new result that is presented here for the first time.
@@
as given by Equation (1.4). We define $v_{-}(x,t)=S(t)u_{-}^{1/\alpha}(x)$ and $v_{+}(x,t)=S(t)u_{+}^{1/\alpha}(x)$. If $v_{0}\in C(\overline{\Omega})$ satisfies $v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)$ in $\overline{\Omega}$, then there exists a unique classical solution $v(x,t)$ to (1.1)-(1.3), such that $v_{-}(x,t)\leq v(x,t)\leq v_{+}(x,t)$ for all $\left(x,t\right)\in\overline{\Omega}\times\left[0,T\right)$.
A careful examination of Theorems 1.1 and 1.3 and their proofs shows that both results play an essential role, even though Theorem 1.1 is a particular case of Theorem 1.3. Theorem 1.1 remains significant because its proof lends itself to straightforward implementation in a programming language through simplified numerical procedures, whereas Theorem 1.3 applies to a more general situation in which the solution is not necessarily separable. As an illustration, we present a well-known result that highlights the distinct roles played by Theorems 1.1 and 1.3.
@@
Therefore, Theorem 1.2 strengthens both the first (Theorem 1.1) and third result (Theorem 1.3). First, it plays a key role in the proof of Theorem 1.3. Second, when combined with Theorem 1.1, it enables the construction of a subsolution and supersolution pair, which in turn makes the implementation of the results in a computer language feasible.
1.7. Organization of the paper
The remainder of this paper is organized as follows. Section 2 presents the Bernoulli localization principle in detail, explaining why it constitutes a new structural framework for this class of equations. Section 3 provides complete proofs of Theorems 1.11.3. In Section 4, we derive the model from conservation laws and energy considerations, following the framework of Gurtin and MacCamy. Section 5 discusses cross-disciplinary relevance, presenting concrete models and verifiable predictions for physics, biology, economics, and engineering. Section 6 provides the complete Python implementation for image denoising and presents a detailed comparison with the Perona–Malik filter, including a transparent explanation of every component of the code. Finally, Section 7 presents conclusions and future research directions.
2. The Bernoulli Localization Principle: A New Structural Framework
In this section we describe in detail the central conceptual contribution of this paper: the Bernoulli localization principle. We explain what it asserts, why it is new, and how it changes the way one can think about the solutions of porous-medium-type diffusion equations with time-dependent growth.
2.1. The principle
Definition 2.1 (Bernoulli localization).
We say that the solutions of the parabolic problem (1.1)–(1.2)–(1.3) are localized by the Bernoulli ODE if there exist functions $v_{-}(x,t)$ and $v_{+}(x,t)$ of the product form
$v_{\pm}(x,t)=S(t)\,u_{\pm}^{1/\alpha}(x),$ (2.1)
where $S(t)$ is the unique positive solution of the Bernoulli-type ODE
$S^{\prime}(t)+S^{\alpha}(t)=\mu(t)\,S(t),\qquad S(0)=\gamma>0,$ (2.2)
and $_{-}, u+u_{+}$ are, respectively, a subsolution and a supersolution of the stationary elliptic problem (1.5) with $u_{-}\leq u_{+}$ in $\overline{\Omega}$, such that every classical solution $v$ of (1.1)–(1.2)–(1.3) with initial data satisfying $v_{-}(x,0)\leq v_{0}(x)\leq v_{+}(x,0)$ is trapped between these barriers:
The power of this definition lies in the explicit, factored form of the barriers. The temporal factor $S(t)$ is universal: it depends only on $\alpha$, $\mu(\cdot)$, and the initial parameter $\gamma$, not on the spatial domain $\Omega$ or the spatial profile $u$. It is given in closed form by Equation (1.4). The spatial factors $u_{\pm}^{1/\alpha}$ are determined by the stationary elliptic problem on $\Omega$—a problem for which existence, uniqueness, and regularity are classical [5, 14].
2.2. Why the principle is new
The novelty of the Bernoulli localization principle can be appreciated by contrasting it with the approaches that have dominated the literature.
  1. 1.
    Classical comparison principles [20, 24] establish that if a subsolution lies below a supersolution at $t=0$ and on the parabolic boundary, then the same ordering persists for all time. These principles are extremely powerful but remain qualitative: they do not specify the functional form of the sub- and supersolutions, nor do they relate the temporal and spatial dependences explicitly. The Bernoulli localization principle, by contrast, produces barriers of a specific, analytically tractable product form, thereby converting a qualitative ordering into a quantitative structural description.
  2. 2.
    Energy and variational methods [24, 12] establish existence via minimization of suitable energy functionals. These methods yield weak solutions and, through regularity bootstrapping, classical solutions. However, the solutions so obtained remain implicit; the energy functional tells us that a minimizer exists, but not what it looks like. The Bernoulli localization principle, in contrast, tells us precisely what the solution looks like: it is the product of a temporal Bernoulli factor and a spatial elliptic profile.
  3. 3.
    Self-similar and asymptotic analyses [1, 25] provide information about the long-time behavior of solutions, typically showing convergence to a self-similar profile or a steady state. The Bernoulli localization principle subsumes these results: the long-time behavior of $S(t)$ is immediately readable from the explicit formula (1.4), yielding the asymptotic profile without further analysis.
2.3. Structural consequences
The Bernoulli localization principle has several immediate structural consequences.
Complete structural description. Every solution of (1.1)–(1.2)–(1.3) lying in the ordered interval $v_{-},v_{+}]$ is sandwiched between two explicitly known functions. In the separable case, the solution is such a product, and the description is exact rather than merely bounding.
Temporal–spatial decoupling. The full parabolic PDE is replaced by two independent, simpler problems: a scalar Bernoulli ODE for the temporal factor and a stationary elliptic PDE for the spatial profile. Each can be analyzed and solved independently, and their combination yields the solution of the original problem.
Explicit parameter dependence. The dependence of the solution on the parameters $\alpha$, $\gamma$, $\mu(\cdot)$, and $\Omega$ is transparently readable from the formula (1.4) and the elliptic problem (1.5). This facilitates sensitivity analysis, parameter fitting in applications, and the design of numerical experiments.
Algorithmic accessibility. The localization principle translates directly into an algorithm: (i) solve the elliptic problem (1.5) by a standard numerical method (finite elements, finite differences, or the iterative scheme in Theorem 1.1); (ii) evaluate $S(t)$ from the closed-form expression (1.4); (iii) form the product $v(x,t)=S(t)\,u^{1/\alpha}(x)$. No parabolic solver is needed for the separable case.